OpenStudy (anonymous):

How are the formulas for the volume of a cylinder, pyramid, and cone derived?

OpenStudy (anonymous):

@TheSmartOne @sleepyjess

OpenStudy (amistre64):

using calculus

OpenStudy (anonymous):

I guess

OpenStudy (anonymous):

But my subject is geo

TheSmartOne (thesmartone):

And calculus is way beyond my level .-.

OpenStudy (igreen):

Well, the formula for the volume of a cone is 1/3 the volume of the cylinder.

TheSmartOne (thesmartone):

@iGreen They want to know how to derive it. :P

OpenStudy (igreen):

Lol, I know.. :P

OpenStudy (igreen):

If anybody can do it @amistre64 can.

OpenStudy (anonymous):

I don't need it in special formula but it being written out would be nice

OpenStudy (anonymous):

@iGreen i hope so.

TheSmartOne (thesmartone):

This explains it nicely too: http://www.shmoop.com/basic-geometry/volume-pyramids-cones.html

OpenStudy (amistre64):



OpenStudy (amistre64):



OpenStudy (amistre64):

\[\int_{0}^{h}\frac{\pi}{m^2}u^2du\] \[\frac{\pi}{3m^2}h^3\] \[\frac{\pi}{3}\frac{hhh}{mm}~:~r=h/m\] \[\frac{\pi}{3}r^2h\]

OpenStudy (amistre64):

tada!!! lol